# Semantic Data Augmentation Enhanced Invariant Risk Minimization for Medical Image Domain Generalization

### v1

@@ IV-A Datasets
The proposed method is evaluated using the GRDBench dataset [10]. GRDBench, introduced by Che et al. [10], serves as a benchmark for domain generalization and consists of eight popular datasets along with two evaluation settings. Due to redistribution restrictions, our experiments focus on four of these datasets: IDRID [32], APTOS [33], RLDR [34], and DEEPDR [35]. The datasets used are APTOS (3662 samples), DEEPDR (1999 samples), IDRID (516 samples), and RLDR (1593 samples). Figure 3 shows samples from each dataset, highlighting the diversity and complexity of the images. Although our method was specifically evaluated on these datasets, its principles of semantic data augmentation and invariant risk minimization suggest potential applicability to other medical image analysis tasks.
Figure 3: Samples of different center dataset of GRDBench [10].
IV-B Implementation Details
For the hyperparameters, the learning rate is set to 0.0001 and the batch size to 256 for all experiments, except for GDRNet [10], where a batch size of 128 is used. This is due to GDRNet’s augmentation strategy, which feeds both the original and augmented images in each iteration, effectively doubling the input size. Hence, a batch size of 128 in GDRNet corresponds to processing 256 images per iteration.
The best learning rate and batch size were selected from { 0.001, 0.0001, 0.0005, 0.00005, 0.00001} and {16, 32, 64, 128, 256}, respectively, based on the validation set. The Adam optimizer [36] is used with a weight decay of 0.00001. All experiments were run for 100 epochs, with the learning rate decayed by a factor of 0.1 at epochs 50 and 75. The backbone network used for all experiments is ResNet-18 [37], with an input size of $224\times 224$ for the GRDBench dataset. PyTorch is used for the deep learning framework, and all experiments were conducted on an NVIDIA RTX 3090 GPU.
IV-C Experiment on GRDBench
We conducted an extensive evaluation to compare the performance of our proposed method with several state-of-the-art approaches on the GRDBench dataset [10]. The comparison includes a vanilla baseline method (ERM)[29], along with advanced techniques from various domains such as ophthalmic disease diagnosis (Fishr, GDRNet)[31, 28], domain generalization (DGT) methods like IRM [16] and VREx [28], and feature representation learning methods (GREEN, CABNet)[9, 30]. These methods were selected based on their relevance to domain generalization and their adaptability to medical image tasks, with minimal to no modifications required for application to our dataset. For all methods, we used the standard domain generalization augmentation pipeline[10] except DRGen, which employed a default augmentation strategy [10].
@@ IV-C Experiment on GRDBench
Results: The results, shown in table I, demonstrate that our method outperforms all other approaches across multiple metrics (AUC, ACC, F1) on most test domains, particularly on DEEPDR, APTOS, and RLDR. Notably, our method achieved the highest AUC in these domains, surpassing other methods by a significant margin. While some methods like Fishr [31] and GDRNet [10] performed well in certain tests, our method consistently outperformed them in key areas, such as the average F1 score and AUC across the four test domains. The results also reveal that domain generalization methods (such as IRM [16] and VREx [28]) show significant improvements over the baseline ERM, but our approach further refines generalization by learning more robust features that preserve diagnostic patterns while increasing intra-class variation. Furthermore, compared to the feature representation learning methods like GREEN [9] and CABNet [30], our method shows notable improvements in both accuracy and robustness, making it a superior choice for domain generalization in medical image tasks. These results underscore the effectiveness of our approach in addressing domain shifts and improving generalization performance across multiple medical imaging datasets.
IV-D Analysis via OTDD
Figure 4: Heatmap of the Optimal Transport Dataset Distance (OTDD)[39] for different domain pairs in the GRDBench dataset [10].
@@ IV-E Ablation Study
Figure 5: Visualizing deep features on GRDBench dataset with two domains using UMAP[40]. For easier demonstration of the effectiveness of the proposed method, we visualized data from only two domains, although the original setting included four domains. As shown in the green box in fig. 5(b), the proposed method consistently enhances features towards the target domain. In contrast, the VIRM method (as shown in fig. 5(a)) produces many random and meaningless directions, especially in the small-batch data scenario, leading to less stable feature representations.
The goal of the direction selector is to make domain A move in the direction of domain B, the This can be achieved by optimizing the direction selector by calculating the difference between the augmented features and other domain features. We consider two potential methods for direction selectors, including maximum mean difference (MMD [24]), covariance of the different source domains (Cov [38]), and domain discriminator [41]. Additionally, we consider potential two different settings for the direction selector: Soft and Hard. Soft indicates that $\mathbb{d}\in[0,1]$ and Hard indicates that $\mathbb{d}\in\{0,1\}$. Table II shows the results of the ablation study, where we compare the performance of our method with different direction selectors and settings.
To further evaluate the proposed method’s ability to reduce domain discrepancies, we employed the Optimal Transport Dataset Distance (OTDD)[39]. OTDD measures the distance between two datasets using optimal transport, where smaller values indicate reduced domain differences. Alvarezet al. [39] demonstrated a linear relationship between OTDD and transferability performance, showing a strong and significant correlation, which highlights OTDD’s predictive power for transferability across datasets.
We innovatively apply OTDD to assess the generalization between two distinct domains within the same dataset. As shown in fig. 3, our method, excluding the GREEN method [9], achieves the lowest average OTDD across domain pairs, indicating superior generalization across diverse domains. The GREEN method [9] performs better in OTDD due to its use of graph convolutional networks (GCNs) to model inter-class dependencies. Since OTDD evaluates label-to-label distances, this results in improved OTDD performance for GREEN. However, our method exhibits significantly better classification performance compared to GREEN, highlighting its practicality in classification tasks.
@@
Figure 6: KDE-Based Visualization of Deep Features Using UMAP[40].
IV-F Visualization
IV-F1 Compare with VIRM
@@
To further emphasize the core advantages of our method, we compared its feature space visualization with that of the VIRM[20] method. As shown in the green box in fig. 5(a), VIRM employs a strategy of randomly selecting augmentation directions, which results in a more dispersed feature space visualization. This suggests that the feature representations learned by VIRM are less stable and compact in the feature space. While this approach may suffice for large datasets that can cover most of the feature space, it becomes less robust for small-batch datasets.
@@
In small-batch settings, the randomness in selecting augmentation directions can lead to meaningless transformations, exacerbating the dispersion of feature space representations. In contrast, as illustrated in the green box in fig. 5(b), our method introduces a domain-covariance-guided augmentation direction selector. This ensures that features are consistently enhanced toward the target domain, resulting in a tighter and more stable feature space overlap. As a result, the feature space visualization of our method is more compact and stable.
@@
Zhu et al. [20] proposed the use of semantic data augmentation to improve IRM-based methods by enhancing feature overlap through semantic diversity, thereby improving model generalization in large-scale datasets. Similarly, our method demonstrates superior performance on small-batch medical datasets by guiding feature enhancement toward the target domain. This ensures more stable and compact feature space representations, which facilitate better generalization.
@@
While fig. 1 provides an idealized model, fig. 5 presents the visualization of our method applied to real-world datasets. These results further validate the effectiveness of our approach and, more specifically, the augmentation direction selector we proposed.
IV-F2 KDE Visualization
@@
Feature support overlap is a crucial determinant of successful generalization[42]. Furthermore, it reflects the extent of dataset diversity shifts along the feature dimensions. To evaluate this overlap, we employed kernel density estimation (KDE)[43], which provides a clear visualization of the feature space overlap. As shown in Figure 6, our method demonstrates notable overlap in the GRDBench dataset, offering an intuitive representation of its effectiveness in reducing domain differences and enhancing generalization.
IV-F3 Feature Visualization
@@
We also visualized the distribution of features from all domains within the same class in the feature space. As illustrated in fig. 7, the features enhanced by our method consistently move toward directions that align with other domains, rather than following random directions. This alignment results in a more compact feature space overlap, which significantly contributes to improving the generalization performance of the model.
@@
Figure 7: UMAP[40]. Visualization of Class 0 Deep Features on the GRDBench Dataset[10] (Original Features Only).

### v2

@@ IV-A Datasets
The proposed method is evaluated using the GRDBench dataset [dgdr2023]. GRDBench, introduced by Che et al. [dgdr2023], serves as a benchmark for domain generalization and consists of eight popular datasets along with two evaluation settings. Due to redistribution restrictions, our experiments focus on four of these datasets: IDRID [IDRID], APTOS [APTOS], RLDR [RLDR], and DEEPDR [DEEPDR]. The datasets used are APTOS (3662 samples), DEEPDR (1999 samples), IDRID (516 samples), and RLDR (1593 samples). Fig.˜3 shows samples from each dataset, highlighting the diversity and complexity of the images. Although our method was specifically evaluated on these datasets, its principles of semantic data augmentation and invariant risk minimization suggest potential applicability to other medical image analysis tasks.
Figure 3: Samples of different center dataset of GRDBench [dgdr2023].
IV-B Evaluation Protocol
To assess the generalization performance of different methods, we follow the leave-one-domain-out cross-validation strategy recommended by GRDBench [dgdr2023]. Specifically, in each experimental run, we designate one of the domains (APTOS, DEEPDR, IDRID, or RLDR) as the test domain, using the remaining three domains for training and validation. This ensures that the model is evaluated on a domain it has never seen during training, simulating real-world domain shifts and enabling robust comparison.
IV-C Implementation Details
For the hyperparameters, the learning rate is set to 0.0001 and the batch size to 256 for all experiments, except for GDRNet [dgdr2023], where the learning rate of 0.001 is used. The best learning rate and batch size were selected from { 0.001, 0.0001, 0.0005, 0.00005, 0.00001} and {16, 32, 64, 128, 256}, respectively, based on the validation set. The Adam optimizer [kingma2014adam] is used with a weight decay of 0.00001. All experiments were run for 100 epochs, with the learning rate decayed by a factor of 0.1 at epochs 50 and 75. The backbone network used for all experiments is ResNet-18 [he2016deep], with an input size of $224\times 224$ for the GRDBench dataset. PyTorch is used for the deep learning framework, and all experiments were conducted on an NVIDIA RTX 3090 GPU.
IV-D Experiment on GRDBench
We conducted an extensive evaluation to compare the performance of our proposed method with several state-of-the-art approaches on the GRDBench dataset [dgdr2023]. The comparison includes a vanilla baseline method (ERM)[vapnik1999nature], along with advanced techniques from various domains such as ophthalmic disease diagnosis (Fishr, GDRNet)[rame2021ishr, krueger2021out], domain generalization (DGT) methods like IRM [arjovsky2019invariant] and VREx [krueger2021out], and feature representation learning methods (GREEN, CABNet)[liu2020green, he2020cabnet]. These methods were selected based on their relevance to domain generalization and their adaptability to medical image tasks, with minimal to no modifications required for application to our dataset. For all methods, we used the standard domain generalization augmentation pipeline[dgdr2023] except DRGen, which employed a default augmentation strategy [dgdr2023].
@@ IV-D Experiment on GRDBench
Results: The results, shown in Table˜I, demonstrate that our method outperforms all other approaches across multiple metrics (AUC, ACC, F1) on most test domains, particularly on DEEPDR, APTOS, and RLDR. Notably, our method achieved the highest AUC in these domains, surpassing other methods by a significant margin. While some methods like Fishr [rame2021ishr] and GDRNet [dgdr2023] performed well in certain tests, our method consistently outperformed them in key areas, such as the average AUC across the four test domains. The results also reveal that domain generalization methods (such as IRM [arjovsky2019invariant] and VREx [krueger2021out]) show significant improvements over the baseline ERM, but our approach further refines generalization by learning more robust features that preserve diagnostic patterns while increasing intra-class variation. Furthermore, compared to the feature representation learning methods like GREEN [liu2020green] and CABNet [he2020cabnet], our method shows notable improvements in both accuracy and robustness, making it a superior choice for domain generalization in medical image tasks. These results underscore the effectiveness of our approach in addressing domain shifts and improving generalization performance across multiple medical imaging datasets.
Figure 4: Heatmap of the Optimal Transport Dataset Distance (OTDD)[alvarez2020geometric] for different domain pairs in the GRDBench dataset [dgdr2023].
@@ IV-E Ablation Study
Figure 5: Visualizing deep features on GRDBench dataset with two domains using UMAP[mcinnes2018umap]. For easier demonstration of the effectiveness of the proposed method, we visualized data from only two domains, although the original setting included four domains. As shown in the green box in Figure˜5(b), the proposed method consistently enhances features towards the target domain. In contrast, the VIRM method (as shown in Figure˜5(a)) produces many random and meaningless directions, especially in the small-batch data scenario, leading to less stable feature representations.
@@
Figure 6: KDE-Based Visualization of Deep Features Using UMAP[mcinnes2018umap].
The goal of the direction selector is to make domain A move in the direction of domain B. This can be achieved by optimizing the direction selector by calculating the difference between the augmented features and other domain features. We consider two potential methods for direction selectors, including maximum mean difference (MMD [li2018domain]), covariance of the different source domains (Cov [wang2024inter]), and domain discriminator [bai2021decaug]. Additionally, we consider potential two different settings for the direction selector: Soft and Hard. Soft indicates that $\mathbb{d}\in[0,1]$ and Hard indicates that $\mathbb{d}\in\{0,1\}$. Table˜II shows the results of the ablation study, where we compare the performance of our method with different direction selectors and settings.
IV-F Analysis via OTDD
To further evaluate the proposed method’s ability to reduce domain discrepancies, we employed the Optimal Transport Dataset Distance (OTDD)[alvarez2020geometric]. OTDD measures the distance between two datasets using optimal transport, where smaller values indicate reduced domain differences. Alvarezet al. [alvarez2020geometric] demonstrated a linear relationship between OTDD and transferability performance, showing a strong and significant correlation, which highlights OTDD’s predictive power for transferability across datasets.
We innovatively apply OTDD to assess the generalization between two distinct domains within the same dataset. As shown in Fig.˜4, our method, excluding the GREEN method [liu2020green], achieves the lowest average OTDD across domain pairs, indicating superior generalization across diverse domains. The GREEN method [liu2020green] performs better in OTDD due to its use of graph convolutional networks (GCNs) to model inter-class dependencies. Since OTDD evaluates label-to-label distances, this results in improved OTDD performance for GREEN. However, our method exhibits significantly better classification performance compared to GREEN, highlighting its practicality in classification tasks.
IV-G Visualization
IV-G1 Compare with VIRM
@@
To further emphasize the core advantages of our method, we compared its feature space visualization with that of the VIRM[zhu2024enlarging] method. As shown in the green box in Figure˜5(a), VIRM employs a strategy of randomly selecting augmentation directions, which results in a more dispersed feature space visualization. This suggests that the feature representations learned by VIRM are less stable and compact in the feature space. While this approach may suffice for large datasets that can cover most of the feature space, it becomes less robust for small-batch datasets.
@@
In small-batch settings, the randomness in selecting augmentation directions can lead to meaningless transformations, exacerbating the dispersion of feature space representations. In contrast, as illustrated in the green box in Figure˜5(b), our method introduces a domain-covariance-guided augmentation direction selector. This ensures that features are consistently enhanced toward the target domain, resulting in a tighter and more stable feature space overlap. As a result, the feature space visualization of our method is more compact and stable.
@@
Zhu et al. [zhu2024enlarging] proposed the use of semantic data augmentation to improve IRM-based methods by enhancing feature overlap through semantic diversity, thereby improving model generalization in large-scale datasets. Similarly, our method demonstrates superior performance on small-batch medical datasets by guiding feature enhancement toward the target domain. This ensures more stable and compact feature space representations, which facilitate better generalization.
@@
While Fig.˜1 provides an idealized model, Fig.˜5 presents the visualization of our method applied to real-world datasets. These results further validate the effectiveness of our approach and, more specifically, the augmentation direction selector we proposed.
IV-G2 KDE Visualization
@@
Feature support overlap is a crucial determinant of successful generalization[ahuja2021invariance]. Furthermore, it reflects the extent of dataset diversity shifts along the feature dimensions. To evaluate this overlap, we employed kernel density estimation (KDE)[wkeglarczyk2018kernel], which provides a clear visualization of the feature space overlap. As shown in Fig.˜6, our method demonstrates notable overlap in the GRDBench dataset, offering an intuitive representation of its effectiveness in reducing domain differences and enhancing generalization.
IV-G3 Feature Visualization
@@
We also visualized the distribution of features from all domains within the same class in the feature space. As illustrated in Fig.˜7, the features enhanced by our method consistently move toward directions that align with other domains, rather than following random directions. This alignment results in a more compact feature space overlap, which significantly contributes to improving the generalization performance of the model.
@@
Figure 7: UMAP[mcinnes2018umap]. Visualization of Class 0 Deep Features on the GRDBench Dataset[dgdr2023] (Original Features Only).